Please help me solve this equation:
4x + 8y = 20
-4x + 2y = -30

Answer by algebrahouse.com(1659)

4x + 8y = 20
-4x + 2y = -30
---------------
10y = -10 {added the two equations together}
y = -1 {divided each side by 10}

4x + 8y = 20 {top original equation}
4x + 8(-1) = 20 {substituted -1, for y, into top original equation}
4x - 8 = 20 {multiplied}
4x = 28 {added 8 to each side}
x = 7 {divided each side by 4}

x = 7 and y = -1

For more help from me, visit: www.algebrahouse.com